OpenStudy (anonymous):

2. Set up the following word problem as a proportion and solve. Then write your answer as a complete sentence. A health department estimates that 8 vials of malaria serum will treat 100 people. At this rate, how many vials are required to treat 175 people? @mathmale

OpenStudy (mathmale):

The question can be reworded a bit: "If 8 vials of serum will treat 100 people, how many vials are necessary to treat 175 people?"

OpenStudy (mathmale):

Obviously, if you have more people, you need more medication / serum.

OpenStudy (mathmale):

Let x represent the number of vials needed to treat 175 people. then \[\frac{ x }{ 8 }\] is the ratio / proportion of the numbers of vials.

OpenStudy (mathmale):

Write the proportion of the # of people: 175/100.

OpenStudy (anonymous):

wouldnt there be 7 people

OpenStudy (mathmale):

Equate the two ratios. Solve the equation for x.

OpenStudy (anonymous):

vials i mean

OpenStudy (anonymous):

im confused

OpenStudy (mathmale):

Vials are glass containers, like tubes with their bottom ends closed.

OpenStudy (anonymous):

i know that i mean the problem

OpenStudy (mathmale):

The pertinent equation is \[\frac{ x }{ 8 }=\frac{ 175 }{ 100 }\] Using cross multiplication, solve this for x. That will tell you how many 'vials' of medication are needed to treat 175 people. Sorry, Purdy, but I have to get off the 'Net now; I'll be back later this evening.